3) Emploved Women If 60% of all women are employed outside the home, find the probablity that in a sample of 9 women,
a. Exactly 6 are employed.
b. At least 4 are employed.
c. Fewer than 3 are employed outside the home.

Answer by ikleyn(52797)

.

These are standard problems on binomial distribution.


(a)   P = P(n=9; k=6; p=0.6).


(b)   P = P(n=9; k >= 4; p=0.6).


(c)   P = P(n=9; k < 3; p=0.6).


Go to web-site https://stattrek.com/online-calculator/binomial and use online (free of charge) calculator there.
Its interface is very simple, while its output is very convenient, so even a beginner student can easily work with it.
It facilitates using  many-addend formula, so calculations are really easy and fast.


Alternatively, you may use standard function binomcfd on your regular TI-83/84 calculator.

About this function and how to use it read from this web-page https://www.statology.org/binompdf-vs-binomcdf/

------------------

To learn on solving typical problems on calculations binomial probabilities,  look into the lessons
    - Elementary operations on sets help solving Probability problems
    - Elementary operations on sets help solving Probability problems - REVISITED
in this site.

You will find there many other similar  (and different)  solved problems.


Learn the subject from there.